OpenStudy (anonymous):

a line passes through a point (-2,3).find an equation of this line if it is 3x+2y=5

OpenStudy (anonymous):

a line passes through a point (-2,3). find an equation of this line if it is ___?____ 3x+2y=5 parallel or perpendicular?

OpenStudy (anonymous):

parallel

OpenStudy (anonymous):

The slope of 3x+2y=5 is m=-3/2 and so the is the slope of the parallel line. use point-slope formula y-y1=m(x-x1) y-3=(-3/2)(x+2) y=(-3/2)x-3+3 y=(-3/2)x

OpenStudy (anonymous):

thanks that's the answer

OpenStudy (anonymous):

ur welcome :})

OpenStudy (anonymous):